# V2Xum-LLM: Cross-Modal Video Summarization with Temporal Prompt Instruction Tuning

### v1

@@ 4.3 Temporal-Aware Decoding
where the temporal token $t_{j}$ is bind with the words or sentences that consist of the summarization of visual content and can be further extracted from $A^{t}$ to get the video-to-video summarization $A^{v}$. This is called temporal-aware decoding. An example is shown in Fig. 3. When decoding, the temporal tokens can be decoded together with the text summaries and represent the temporal position where the described visual content occurred in the input video.
4.4 Task-Controllable Video Summarization
As mentioned before, the types of summarization can be controlled by the instructions $I^{x}$, and we will further introduce how to utilize the instruction following the capability of LLM to realize the modality controllable video summarization.
We use instructions like “Please generate a/BOTH VIDEO/TEXT summarization for this video.” as the instructions and tokenize them with the same text tokenizer that tokenizes the temporal prompts.
During the instruction tuning, we freeze the parameters of the vision encoder and update the vision adapter, an MLP that aligns the visual semantic with the input token space of the LLM, and the language decoder to force it to generate temporal indexes.
5 Experiments
Table 2: Comparison Results on the VideoXum dataset. “TSH-Free” means the model is task-specific-head-free; “B-4” denotes the BLEU-4; “M” denotes the METEOR metric; “R-L” refers to the ROUGE-L metric; “C” represents the CIDEr metric for consensus-based evaluation. For all evaluation metrics, higher scores indicate better performance.
@@ B-4 M R-L C F1 Spearman Kendall FCLIP Cross-FCLIP
DENSE [31] 1.6 8.9 - - - - - - -
DVC-D-A [34] 1.7 9.3 - - - - - - -
Bi-LSTM+TempoAttn [93] 2.1 10.0 - - - - - - -
Masked Transformer [93] 2.8 11.1 - - - - - - -
Support-Set [52] 1.5 6.9 17.8 3.2 - - - - -
Frozen-BLIP [33] 0.0 0.4 1.4 0.0 16.1 0.011 0.008 - -
Vid2Seq-HCY [81] 2.3 8.2 19.0 7.6 24.2 - - 0.888 0.214
Vid2Seq-HC [81] 2.7 8.5 19.8 8.4 24.5 - - 0.892 0.217
Vid2Seq-HCV [81] 2.7 8.4 19.8 8.3 25.1 - - 0.899 0.200
VSUM-BLIP [37] - - - - 21.7 0.207 0.131 - -
TSUM-BLIP [37] 5.6 11.8 24.9 20.9 - - - - -
VTSUM-BLIP [37] 5.8 12.2 25.1 23.1 23.5 0.258 0.196 0.894 0.247
@@
Table 3: Comparison results on the TVSum and SumMe datasets.
Method TVSum SumMe
Spearman Kendall Spearman Kendall
dppLSTM [87] 0.055 0.042 - -
DSN [91] 0.020 0.026 - -
Sumgraph [51] 0.138 0.094 - -
CLIP-it [48] 0.147 0.108 0.120 0.109
TL;DW [47] 0.167 0.143 0.128 0.111
iPTNet [27] 0.174 0.148 0.131 0.114
A2Summ [17] 0.178 0.150 0.143 0.121
Standard ranker [56] 0.230 0.176 0.014 0.011
VSUM-BLIP [37] 0.261 0.200 0.365 0.268
V2Xum-LLaMA 0.293 0.222 0.378 0.296
5.1 Baseline Models
Our V2Xum-LLaMA model, including both 7B and 13B versions, is benchmarked against various models adept at video-to-video (V2V) summarization, video-to-text (V2T) summarization, or both tasks on the VideoXum dataset. Baseline models include LLM-based approaches such as Frozen-BLIP [33], VSUM-BLIP [37], TSUM-BLIP [37], and VTSUM-BLIP [37]. We compare with task-specific-head-free (TSH-Free) models like DENSE [31], DVC-D-A [34], Bi-LSTM+TempoAttn [93], Masked Transformer [93], and Support-Set [52], which do not rely on regression-based timestamp prediction with extra task-specific heads. Additionally, on the classical TVSum [61] and SumMe [14] datasets, we compare our 7B version V2Xum-LLaMA with the following V2V summarization methods: dppLSTM [87], DSN [91], Sumgraph [51], CLIP-it [48], TL;DW [47], iPTNet [27], A2Summ [17], Standard ranker [56], and VSUM-BLIP [37]. We also evaluated Vid2Seq [81] models, initially pre-trained on the HowTo100M [45], VidChapter-7M [80], YouCook2 [92], and ViTT [23] datasets, and subsequently fine-tuned on the VideoXum dataset to establish baseline models for our experiments.
5.2 Evaluation Metrics
We introduce new CLIP-based evaluation metrics for V2V and V2VT summarization evaluation. While the F1 score is a common metric for V2V summarization tasks, the process of video summarization annotation is highly subjective, leading to considerable variance among human annotators [49]. The traditional F1 score, which compares predicted video frames directly with the ground truth, fails to account for semantically similar frames that are close in time but not exactly matching, thus potentially undervaluing accurate summaries. To mitigate this, we introduce the $F_{CLIP}$ metric for V2V summarization evaluations, which is designed to recognize and reward semantic similarities between predicted and ground truth frames even when they are not identical. And we also propose the $Cross-F_{CLIP}$ metric for the V2VT summarization tasks. Unlike the VT-CLIPScore metric used in VideoXum—which calculates the average cross-modal CLIP score as an indicator of semantic alignment between predicted video and text summaries—our $Cross-F_{CLIP}$ calculates the average $F_{CLIP}$ scores between the predicted video summaries and the ground truth text summaries, as well as between the predicted text summaries and the ground truth video summaries. This approach aims to provide a more nuanced evaluation of summarization tasks by acknowledging the importance of semantic content alignment across modalities. For a reference video summary $v$ and predicted video summary $\hat{v}$, the recall, precision, and F1 scores are:
@@ 5.3 Implementation Details
We use CLIP ViT-L/14@336 as the vision encoder and Vicuna-v1.5-7B/13B as the large language model. The vision adapter is pre-trained by vision-text alignment with image-text pairs from LCS-558K [39] dataset. After being processed by the Vicuna-v1.5 tokenizer, each temporal prompt takes four tokens. To facilitate model training, we normalize the video length into 100 via downsampling. For each setting, we train the V2Xum-LLaMA for 5 epochs with a learning rate of 1e-4 on 8$\times$×NVIDIA A100 GPUs.
5.4 Quantitative Results
@@
To verify the effectiveness of our proposed method, we conduct experiments on multiple video summarization benchmarks, including cross-modal video summarization and video summarization tasks. For cross-modal video summarization, we evaluate our method using the VideoXum dataset [37] along with our newly proposed V2Xum dataset. In the realm of V2V summarization, we utilized TVSum [61] and SumMe [14] as benchmarks for our evaluation.
5.4.1 Cross-Modal Video Summarization
@@
The experimental results of VideoXum are shown in Table 2. It can be summarized from the table that our proposed V2Xum-LLaMA model outperforms all the baseline models. Specifically, in the task of V2V summarization, V2Xum-LLaMA achieves a 8.1% higher F1-Score than VTSUM-BLIP, alongside significant improvements in both Spearman and Kendall correlation metrics. In addition, V2Xum-LLaMA-13B achieves higher evaluation scores on the V2V summarization task than V2Xum-LLaMA-7B. On the contrary, V2Xum-LLaMA-7B performs better on V2T summarization. We attribute this result to the increased complexity of the V2T summarization compared to V2V summarization. Additionally, the VideoXum training set comprises only 8,000 examples, a quantity insufficient for effectively training models with large language decoders.
@@
We present the performance of various models on our newly proposed Instruct-V2Xum dataset, with the results detailed in Table 4. The results indicate that the models are well-adapted to the dataset, exhibiting sound performance. Moreover, V2Xum-LLaMA-13B outshines V2Xum-LLaMA-7B in V2T summarization, which we believe is due to the larger language models benefiting from the increased volume of training data.
Table 4: Comparison results on the Instruct-V2Xum test set.
Method V2T V2V V2TV
B-4 M R-L C F1 FCLIP Cross-FCLIP
Vid2Seq-HC [81] 3.8 6.1 22.6 0.4 23.0 80.5 16.1
Vid2Seq-HCY [81] 3.7 6.2 22.4 0.5 24.7 81.3 16.0
Vid2Seq-HCV [81] 3.6 6.2 22.5 0.4 25.1 81.5 16.3
V2Xum-LLaMA-7B 6.8 15.8 26.9 0.9 31.7 95.5 23.1
V2Xum-LLaMA-13B 6.7 15.8 27.0 0.8 31.3 95.3 23.0
5.4.2 Video-to-Video Summarization
@@
We evaluate the capability of V2Xum-LLaMA on VideoXum, TVSum, SumMe, and Instruct-V2Xum datasets. The results are shown in Table 2 and Table 3. It can be concluded from the empirical results that the unified video summarization using the language decoders in V2Xum-LLaMA can effectively perform traditional V2V summarization tasks. V2Xum-LLaMA outperforms all the previous methods that relied on task-specific regression heads for generating video summaries. This result indicates that LLMs with temporal prompts are capable of performing fine-grained video temporal understanding. The results presented in Table 4 affirm the validity of our proposed V2Xum dataset. It demonstrates that the model can properly fit the data.
5.5 Ablation Study
To further evaluate the efficacy of our proposed V2Xum-LLaMA framework and to underscore the importance of augmenting the training dataset, we conduct an ablation study on V2V and V2VT summarization tasks. First, we explore the performance enhancements attributed to the introduction of the temporal prompt mechanism, with findings detailed in Table 5. The data reveal that integrating the temporal prompt into the V2Xum-LLaMA framework yields superior results in both V2V and V2T summarization tasks compared to models lacking this feature, which indicates the temporal prompt benefits language models for fine-grained dense video temporal understanding. Second, to analyze the impact of concurrently generating video and text summaries on the performance of V2V and V2T summarization tasks, we conduct ablation studies. The results, as illustrated in Table 5, suggest that the simultaneous generation of video and text summaries enhances the performance in the V2V summarization task. Third, we also compare the performance of models training with more data (Instruct-V2Xum + VideoXum) and the models finetuned without more data. The results indicate that instruction data can effectively improve models’ performance, especially for the V2T summarization task. Furthermore, we conduct a comparison between models equipped with a pre-trained vision adapter and those without to highlight the critical importance of pre-training the vision adapter in enhancing performance. The results indicate that a well-pre-trained visual prompt adapter enhances the models’ performance for V2V and V2T tasks. Finally, we provide a comparison of parameter-efficient finetuning (PEFT) and full parameter fine-tuning. The results are presented in Table 5. The results demonstrate that full parameter fine-tuning outperforms PEFT in performance.
@@ 5.5 Ablation Study
6 Discussions on Video Summarization: Data, Method, and Evaluation
@@
In this section, we discuss the existing vision summarization tasks from the perspective of data, method, and evaluation. As mentioned before, current existing video summarization datasets, including V2V and V2VT video summarization datasets, contain few training examples and cannot support training large-scale deep neural networks. The TVSum dataset comprises merely 50 YouTube videos; the SumMe dataset includes only 25 personal videos sourced from YouTube, while the QFVS dataset provides 135 video-query training samples. Additionally, VideoXum, the first cross-modal video summarization dataset, provides only 8k training examples. Our experimental results reveal that it is insufficient for training 13B models for the V2T summarization task. This issue is one of the significant drawbacks of current existing datasets. To address this problem, we collect more videos from YouTube and generate the corresponding cross-modal summaries for the videos using GPT-4 and propose a large-scale cross-modal video summarization dataset.
@@
A conventional approach to V2V summarization involves training a regression head to assign an importance score to each frame, ranking frames based on these scores, and selecting the top Kframes to evaluate performance using metrics like the F1 score or Kendall/Spearman correlation against the ground truth. However, as demand for cross-modal video summarization grows, this method is inadequate for multimodal video summarization. Recent studies have begun to explore the use of language models for generating temporal and spatial references in videos [22, 35, 55], demonstrating the viability of using language models to generate the video interval indexes. In addition, leveraging large language models’ text decoders for both V2V and V2T summarization tasks is a logical step. Therefore, in this study, we investigate how to prompt large VLMs to understand the fine-grained video content along with temporal information, how to maximally leverage the powerful capability of content understanding and reasoning of large language models, and how to achieve task controllability via natural language instructions. To that end, we first propose the temporal prompt mechanism, and then design the visual encoder that takes the interleaved video frames, temporal, and language as input. Instead of using the regression head to perform V2V summarization, we unify different video summarization tasks into one language decoder.
@@
Evaluating video summaries poses a significant challenge, primarily because the criteria for quality are inherently subjective, vary across different viewers and even fluctuate over time. This subjectivity, coupled with the limited availability of evaluation videos and annotations, further exacerbates the ambiguity in assessing video summary quality [49]. The traditional F1 score, designed to directly compare predicted video frames with ground truth, ignores the nuances of semantically similar frames that, while temporally proximate, do not exactly match. This oversight can lead to the undervaluation of otherwise accurate summaries. To address this, we design new CLIP-based F scores for evaluating V2V and V2VT summarization tasks. Unlike the traditional F1 score, which relies on exact matches for its precision and recall calculations, $F_{CLIP}$ evaluates the V2V summarization from the perspective of semantic similarity. This approach allows for a more comprehensive yet meaningful evaluation that recognizes the importance of semantic accuracy over mere frame-by-frame accuracy. To evaluate the alignment of the generated video and text summaries, VideoXum employs an approach that calculates the average CLIP score across video frames and text summaries. This method involves computing vectors for all frames and sentences, applying mean pooling to these vectors to represent the video and text summaries, and then calculating the cosine similarity between them. However, this approach evaluates the video and text summaries as unified entities and thus neglects the detailed alignment at the sentence level with specific video frames. In contrast, our proposed $Cross-F_{CLIP}$ adopts a greedy matching strategy to optimize the similarity score between individual video frames and corresponding sentences, ensuring a more granular and accurate alignment.
7 Conclusion

### v2

@@ Introduction
where the temporal token $t_{j}$ is bind with the words or sentences that consist of the summarization of visual content and can be further extracted from $A^{t}$ to get the V2V summarization $A^{v}$. This is called temporal-aware decoding. An example is shown in Fig. 2. When decoding, the temporal tokens can be decoded together with the text summaries and represent the temporal position where the described visual content occurred in the input video.
Task-Controllable Video Summarization Training
As mentioned before, the types of summarization can be controlled by the task instructions $I^{x}$. Specifically, we use task prompts like “Please generate a BOTH/VIDEO/TEXT summarization for this video.” to instruct models to generate the corresponding video summaries. We then train the model end-to-end using negative log-likelihood loss:
$\mathcal{L}=-\sum_{\mathcal{D}}\sum_{i=1}^{N}\log p(A^{x}_{i}\mid\overline{S},A^{x}_{\leq i-1}).$ (4)
where $\mathcal{D}$ denotes the training samples in the dataset, and $N$ is the length of the generated video and text summaries. During the training, we freeze the parameters of the vision encoder and update the vision adapter, and the language decoder to train the model to generate video and text summaries.
Experiments
Baseline Models
We benchmarked our V2Xum-LLaMA model, both 7B and 13B versions, against various models on V2V, V2T, and V2VT summarization tasks using the VideoXum dataset. Baseline models include LLM-based approaches such as Frozen-BLIP (Li et al. 2023), VSUM-BLIP (Lin et al. 2023b), TSUM-BLIP (Lin et al. 2023b), and VTSUM-BLIP (Lin et al. 2023b). We compare with task-specific-head-free (TSH-Free) models like DENSE (Krishna et al. 2017), DVC-D-A (Li et al. 2018), Bi-LSTM+TempoAttn (Zhou et al. 2018), Masked Transformer (Zhou et al. 2018), and Support-Set (Patrick et al. 2020), which do not rely on regression-based timestamp prediction with extra task-specific heads. Additionally, on the classical TVSum (Song et al. 2015) and SumMe (Gygli et al. 2014) datasets, we compare our 7B version V2Xum-LLaMA with the following V2V summarization methods: dppLSTM (Zhang et al. 2016), DSN (Zhou, Qiao, and Xiang 2018), Sumgraph (Park et al. 2020), CLIP-it (Narasimhan et al. 2021), TL;DW (Narasimhan et al. 2022), iPTNet (Jiang and Mu 2022), A2Summ (He et al. 2023a), Standard ranker (Saquil et al. 2021), and VSUM-BLIP (Lin et al. 2023b). We also evaluated Vid2Seq (Yang et al. 2023) models, initially pre-trained on the HowTo100M (Miech et al. 2019), VidChapter-7M (Yang et al. 2024), YouCook2 (Zhou, Xu, and Corso 2018), and ViTT (Huang et al. 2020) datasets, and fine-tuned on VideoXum for fair comparison.
@@ B-4 M R-L C F1 Spearman Kendall FCLIP Cross-FCLIP
DENSE (Krishna et al. 2017) 1.6 8.9 - - - - - - -
DVC-D-A (Li et al. 2018) 1.7 9.3 - - - - - - -
Bi-LSTM+TempoAttn (Zhou et al. 2018) 2.1 10.0 - - - - - - -
Masked Transformer (Zhou et al. 2018) 2.8 11.1 - - - - - - -
Support-Set (Patrick et al. 2020) 1.5 6.9 17.8 3.2 - - - - -
Frozen-BLIP (Li et al. 2023) 0.0 0.4 1.4 0.0 16.1 0.011 0.008 - -
Vid2Seq-HCY (Yang et al. 2023) 2.3 8.2 19.0 7.6 24.2 - - 0.888 0.214
Vid2Seq-HC (Yang et al. 2023) 2.7 8.5 19.8 8.4 24.5 - - 0.892 0.217
Vid2Seq-HCV (Yang et al. 2023) 2.7 8.4 19.8 8.3 25.1 - - 0.899 0.200
VSUM-BLIP (Lin et al. 2023b) - - - - 21.7 0.207 0.131 - -
TSUM-BLIP (Lin et al. 2023b) 5.6 11.8 24.9 20.9 - - - - -
VTSUM-BLIP (Lin et al. 2023b) 5.8 12.2 25.1 23.1 23.5 0.258 0.196 0.894 0.247
@@
Table 2: Comparison Results on the VideoXum dataset. “TSH-Free” indicates the model is task-specific-head-free; “B-4” denotes BLEU-4; “M” denotes METEOR; “R-L” refers to ROUGE-L metric; “C” represents CIDEr. For all metrics, higher scores indicate better performance.
Method V2T V2V V2TV
B-4 M R-L C F1 FCLIP Cross-FCLIP
Vid2Seq-HC (Yang et al. 2023) 3.8 6.1 22.6 0.4 23.0 80.5 16.1
Vid2Seq-HCY (Yang et al. 2023) 3.7 6.2 22.4 0.5 24.7 81.3 16.0
Vid2Seq-HCV (Yang et al. 2023) 3.6 6.2 22.5 0.4 25.1 81.5 16.3
V2Xum-LLaMA-7B 6.8 15.8 26.9 0.9 31.7 95.5 23.1
V2Xum-LLaMA-13B 6.7 15.8 27.0 0.8 31.3 95.3 23.0
Table 3: Comparison results on the Instruct-V2Xum test set.
Evaluation Metrics
We introduce new CLIP-based evaluation metrics for V2V and V2VT summarization evaluation. While the F1 score is a common metric for V2V summarization tasks, the process of video summarization annotation is highly subjective, leading to considerable variance among human annotators (Otani et al. 2019). The traditional F1 score, which compares predicted video frames directly with the ground truth, fails to account for semantically similar frames that are close in time but not exactly matching, thus potentially undervaluing accurate summaries. To mitigate this, we introduce the $F_{CLIP}$ metric for V2V summarization evaluations, which is designed to recognize and reward semantic similarities between predicted and ground truth frames even when they are not identical. And we also propose the $Cross-F_{CLIP}$ metric for the V2VT summarization tasks. Unlike the VT-CLIPScore metric used in VideoXum—which calculates the average cross-modal CLIP score as an indicator of semantic alignment between predicted video and text summaries—our $Cross-F_{CLIP}$ calculates the average $F_{CLIP}$ scores between the predicted video summaries and the ground truth text summaries, as well as between the predicted text summaries and the ground truth video summaries. This approach aims to provide a more nuanced evaluation of summarization tasks by acknowledging the importance of semantic content alignment across modalities. For a reference video summary $v$ and predicted video summary $\hat{v}$, the recall, precision, and F1 scores are:
@@ B-4 M R-L C F1 FCLIP Cross-FCLIP
We use CLIP ViT-L/14@336 as the vision encoder and Vicuna-v1.5-7B/13B as the text decoder. The vision adapter is pre-trained with image-text pairs from LCS-558K (Liu et al. 2023a) dataset. Each tokenized temporal prompt takes four tokens. To facilitate model training, we normalize the video length into 100 via downsampling. For each setting, we train the V2Xum-LLaMA for 5 epochs with a learning rate of 1e-4 on 8$\times$×NVIDIA A100 GPUs.
Quantitative Results
@@
To verify the effectiveness of our method, we conducted experiments on various video summarization benchmarks, including cross-modal and V2V summarization tasks. For cross-modal video summarization, adopt the VideoXum dataset (Lin et al. 2023b) and our proposed V2Xum dataset. For V2V summarization, we used the TVSum (Song et al. 2015) and SumMe (Gygli et al. 2014) benchmarks.
Cross-Modal Video Summarization.
@@
We use the VideoXum dataset to evaluate our model’s capability to cross-model video summarization. The experimental results are shown in Table 2. It can be summarized that our proposed method outperforms all the baseline models. Specifically, in V2V summarization, V2Xum-LLaMA achieves a 8.1% higher F1-Score than VTSUM-BLIP, alongside significant improvements in both Spearman and Kendall correlation metrics. In addition, V2Xum-LLaMA-13B achieves higher evaluation scores on the V2V summarization task than V2Xum-LLaMA-7B. On the contrary, V2Xum-LLaMA-7B performs better on V2T summarization. We attribute this result to the increased complexity of the V2T summarization compared to V2V summarization. Additionally, the VideoXum training set comprises only 8,000 examples, a quantity insufficient for effectively training models with large language decoders.
@@
We also evaluate various models on our newly proposed Instruct-V2Xum dataset, as detailed in Table 3. The results indicate that the models are well-adapted to the dataset, exhibiting sound performance. Moreover, V2Xum-LLaMA-13B outperforms V2Xum-LLaMA-7B in V2T summarization, which we believe is due to the larger language models benefiting from the increased volume of training data.
Video-to-Video Summarization.
@@
We evaluate V2Xum-LLaMA on VideoXum, TVSum, SumMe, and Instruct-V2Xum datasets. The results are shown in Table 2 and Table 4. It can be concluded that the unified video summarization using the language decoders in V2Xum-LLaMA can effectively perform traditional V2V summarization tasks. V2Xum-LLaMA outperforms all the previous methods that relied on task-specific regression heads for generating video summaries. This result indicates that LLMs with temporal prompts are capable of performing fine-grained video temporal understanding. The results presented in Table 3 affirm the validity of our proposed V2Xum dataset. It demonstrates that the model can properly fit the data.
Method TVSum SumMe
Spearman Kendall Spearman Kendall
dppLSTM (Zhang et al. 2016) 0.055 0.042 - -
DSN (Zhou, Qiao, and Xiang 2018) 0.020 0.026 - -
Sumgraph (Park et al. 2020) 0.138 0.094 - -
CLIP-it (Narasimhan et al. 2021) 0.147 0.108 0.120 0.109
TL;DW (Narasimhan et al. 2022) 0.167 0.143 0.128 0.111
iPTNet (Jiang and Mu 2022) 0.174 0.148 0.131 0.114
A2Summ (He et al. 2023a) 0.178 0.150 0.143 0.121
Standard ranker (Saquil et al. 2021) 0.230 0.176 0.014 0.011
VSUM-BLIP (Lin et al. 2023b) 0.261 0.200 0.365 0.268
V2Xum-LLaMA 0.293 0.222 0.378 0.296
Table 4: Comparison results on the TVSum and SumMe datasets.
@@
Ablation Study
To better evaluate the effectiveness of our proposed V2Xum-LLaMA framework and to underscore the importance of augmenting the training dataset, we conduct an ablation study on V2V and V2VT summarization tasks, detailed in Table 5. First, we explore the performance enhancements attributed to the introduction of the temporal prompt mechanism. The results reveal that integrating the temporal prompt into the V2Xum-LLaMA framework yields superior results in both V2V and V2T summarization tasks, indicating that the temporal prompt benefits language models for fine-grained dense video temporal understanding. Second, we evaluated the simultaneous generation of video and text summaries, which enhanced performance in the V2V task. Third, we compared models trained with additional data (Instruct-V2Xum + VideoXum) to those without, showing that instruction data notably boosts performance, particularly in V2T summarization. Additionally, models with a pre-trained vision adapter outperformed those without, emphasizing the importance of pre-training for visual prompts. Finally, we compared parameter-efficient fine-tuning (PEFT) to full parameter fine-tuning, with results favoring the latter in overall performance.
Limitations of Current Video Summarization: Data, Methods, and Evaluation
@@
In this section, we discuss the limitations of existing vision summarization tasks from the perspective of data, method, and evaluation. As mentioned before, current existing video summarization datasets, including V2V and V2VT video summarization datasets, contain few training examples and cannot support training large-scale deep neural networks. The TVSum dataset comprises merely 50 YouTube videos; the SumMe dataset includes only 25 personal videos sourced from YouTube, while the QFVS dataset provides 135 video-query training samples. Additionally, VideoXum, the first cross-modal video summarization dataset, provides only 8k training examples. Our experimental results reveal that it is insufficient for training 13B models for the V2T summarization task. This issue is one of the significant drawbacks of current existing datasets. To address this problem, we collect more videos from YouTube and generate the corresponding cross-modal summaries for the videos using GPT-4 and propose a large-scale cross-modal video summarization dataset.
@@
A conventional approach to V2V summarization involves training a regression head to assign an importance score to each frame, ranking frames based on these scores, and selecting the top Kframes to evaluate performance using metrics like the F1 score or Kendall/Spearman correlation against the ground truth. However, as demand for cross-modal video summarization grows, this method is inadequate for multimodal video summarization. Recent studies have begun to explore the use of language models for generating temporal and spatial references in videos (Huang et al. 2023; Li et al. 2024; Ren et al. 2023), demonstrating the viability of using language models to generate the video interval indexes. In addition, leveraging large language models’ text decoders for both V2V and V2T summarization tasks is a logical step. Therefore, in this study, we investigate how to prompt large VLMs to understand the fine-grained video content along with temporal information, how to maximally leverage the powerful capability of content understanding and reasoning of large language models, and how to achieve task controllability via natural language instructions. To that end, we first propose the temporal prompt mechanism, and then design the visual encoder that takes the interleaved video frames, temporal, and language as input. Instead of using the regression head to perform V2V summarization, we unify different video summarization tasks into one language decoder.
@@
Evaluating video summaries poses a significant challenge, primarily because the criteria for quality are inherently subjective, vary across different viewers and even fluctuate over time. This subjectivity, coupled with the limited availability of evaluation videos and annotations, further exacerbates the ambiguity in assessing video summary quality (Otani et al. 2019). The traditional F1 score, designed to directly compare predicted video frames with ground truth, ignores the nuances of semantically similar frames that, while temporally proximate, do not exactly match. This oversight can lead to the undervaluation of otherwise accurate summaries. To address this, we design new CLIP-based F scores for evaluating V2V and V2VT summarization tasks. Unlike the traditional F1 score, which relies on exact matches for its precision and recall calculations, $F_{CLIP}$ evaluates the V2V summarization from the perspective of semantic similarity. This approach allows for a more comprehensive yet meaningful evaluation that recognizes the importance of semantic accuracy over mere frame-by-frame accuracy. To evaluate the alignment of the generated video and text summaries, VideoXum employs an approach that calculates the average CLIP score across video frames and text summaries. This method involves computing vectors for all frames and sentences, applying mean pooling to these vectors to represent the video and text summaries, and then calculating the cosine similarity between them. However, this approach evaluates the video and text summaries as unified entities and thus neglects the detailed alignment at the sentence level with specific video frames. In contrast, our proposed $Cross-F_{CLIP}$ adopts a greedy matching strategy to optimize the similarity score between individual video frames and corresponding sentences, ensuring a more granular and accurate alignment.
Conclusion